# Statistical-Computational Trade-offs for Recursive Adaptive Partitioning Estimators

### v2

@@ Appendix I Constant Probability Lower Bounds
$h(\mathbf{x};\mathcal{D}_{n})\coloneqq\mathbb{E}_{\theta_{1}}\left\{\hat{f}_{\operatorname{G}}(\mathbf{x};\mathcal{D}_{n},\theta_{1})\mathbb{1}\left\{J(\mathbf{x};\mathcal{D}_{n},\theta_{1})\cap T\neq\emptyset\right\}\right\}.$ (I.9)
A modification of the proof of Lemma C.3 gives
$\begin{split}\mathbb{E}_{\Theta}\left\{R\left(\hat{f}_{\operatorname{G-E}}(-;\mathcal{D}_{n},\Theta)\right)\right\}&\geq\mathbb{E}_{\mathbf{X}}\left\{\frac{\operatorname{Var}\left\{f^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash T}\right\}}{2}-\operatorname{Var}\left\{h(\mathbf{X};\mathcal{D}_{n})~{}|~{}\mathbf{X}_{[d]\backslash T}\right\}\right\}\\ &\geq\frac{\operatorname{Var}\left\{r_{\textsf{MSP}}(\mathbf{X})\right\}}{2}-\mathbb{E}_{\mathbf{X}}\left\{\operatorname{Var}\left\{h(\mathbf{X};\mathcal{D}_{n})~{}|~{}\mathbf{X}_{[d]\backslash T}\right\}\right\}.\end{split}$ (I.10)
@@
$\mathbb{E}_{\mathbf{X}}\left\{\operatorname{Var}\left\{h(\mathbf{X};\mathcal{D}_{n})~{}|~{}\mathbf{X}_{[d]\backslash T}\right\}\right\}\leq\mathbb{E}_{\mathbf{X}}\left\{h(\mathbf{X};\mathcal{D}_{n})^{2}\right\}.$ (I.11)
@@
$\begin{split}\mathbb{E}_{\mathbf{X}}\left\{h(\mathbf{X};\mathcal{D}_{n})^{2}\right\}&\leq\mathbb{E}_{\mathbf{X},\theta_{1}}\left\{\hat{f}_{\operatorname{G}}(\mathbf{X};\mathcal{D}_{n},\theta_{1})^{2}\mathbb{1}\left\{J(\mathbf{x};\mathcal{D}_{n},\theta_{1})\cap T\neq\emptyset\right\}\right\}\\ &\leq M^{2}\mathbb{P}_{\mathbf{X},\theta_{1}}\left\{J(\mathbf{X};\mathcal{D}_{n},\theta_{1})\cap T\neq\emptyset\right\}.\end{split}$ (I.12)
@@
$\mathbb{E}_{\mathcal{D}_{n}}\left\{\mathbb{P}_{\mathbf{X},\theta_{1}}\left\{J(\mathbf{x};\mathcal{D}_{n},\theta_{1})\cap T\neq\emptyset\right\}\right\}\leq\delta,$ (I.13)
@@
$\mathbb{P}_{\mathbf{X},\theta_{1}}\left\{J(\mathbf{x};\mathcal{D}_{n},\theta_{1})\cap T\neq\emptyset\right\}\leq 2\delta.$ (I.14)
@@
$\mathbb{E}_{\Theta}\left\{R\left(\hat{f}_{\operatorname{G-E}}(-;\mathcal{D}_{n},\Theta)\right)\right\}\geq\frac{\operatorname{Var}\left\{r_{\textsf{MSP}}(\mathbf{x})\right\}}{4}.$ (I.15)
@@
Appendix J Comparisons with Neural Networks Trained by SGD
@@
$\left|S_{i}\backslash\cup_{j=1}^{i-1}S_{j}\right|\leq l$ (J.1)
for any $i\in\{1,2,\ldots,r\}$. As can easily be seen, MSP functions are equivalent to leap-$1$ functions, while non-MSP functions are categorized as leap-$2$, leap-$3$, and so on. A Boolean monomial of degree $k$ (i.e., a $k$-parity) is an example of a leap-$k$ function. With this definition, Abbe et al. (2023) (see Conjecture 1 therein) conjectured that the time and sample complexity required to learn a leap-$k$ function using online SGD on a two-layer NN is $\tilde{\Theta}\left(d^{\max\left\{k-1,1\right\}}\right)$.
If this conjecture is true, then combining it with the results in our paper yields rigorous theoretical performance gaps between two-layer NNs trained with online SGD (henceforth denoted as 2LNNOSGD) and greedy regression trees and ensembles: More specifically, for leap-$1$ functions, that is, those satisfying MSP, CART requires a sample complexity of just $O(2^{s}\log d)$ which is much smaller than the $\tilde{\Theta}(d)$ sample complexity required by 2LNNOSGD. On the other hand, for leap-$k$ functions for any $k>1$, CART requires a sample complexity of $\exp(\Omega(d))$, which is much larger than the $\tilde{\Theta}\left(d^{k-1}\right)$ sample complexity required by 2LNNOSGD. Here we note that our lower bound in Theorem 1.1 holds even with noiseless observations and also with constant probability (see Theorem I.1), so that both classes of learning algorithms are put on exactly the same theoretical footing.
@@
The lower bounds portion of Abbe et al. (2023)’s conjecture is suggested by existing Statistical Query (SQ) and Correlational Statistical Query (CSQ) lower bounds. SQ and CSQ are generic models of constrained computation (Kearns, 1998; Feldman et al., 2021). In the SQ model, the analyst has access to a sequence of queries, each of which returns an approximation of the expectation $\mathbb{E}\left\{\psi(\textbf{X},Y)\right\}$ for some real-valued function $\psi$ that may change from query to query. The CSQ model further constrains the queries to be of the form $\mathbb{E}\left\{\phi(\textbf{X})Y\right\}$ for some real-valued function $\phi$. Under the SQ model, it is known that $\Omega\left(d^{k}\right))$ queries are required to learn $k$-parities (Barak et al., 2022), while Abbe et al. (2023) proved a CSQ lower bound of the same order. Now, Abbe et al. (2021) showed that SGD is no more powerful than the SQ model if the stochastic gradients have limited precision, but it is unclear to us whether this holds for the SGD regime in 2LNNOSGD. Assuming the reduction holds, then the SQ and CSQ lower bounds translate to a lower sample complexity bound of $\tilde{\Theta}\left(d^{k-1}\right)$ (Kou et al., 2024).

### v3

@@ Appendix G Stable MSP Properties
$h(\mathbf{x};\mathcal{D}_{n})\coloneqq\mathbb{E}_{\theta_{1}}\left\{\hat{f}_{\operatorname{G}}(\mathbf{x};\mathcal{D}_{n},\theta_{1})\mathbb{1}\left\{J(\mathbf{x};\mathcal{D}_{n},\theta_{1})\cap T\neq\emptyset\right\}\right\}.$ (J.9)
A modification of the proof of Lemma D.3 gives
[…]
@@
[…]
@@
$\begin{split}\mathbb{E}_{\mathbf{X}}\left\{h(\mathbf{X};\mathcal{D}_{n})^{2}\right\}&\leq\mathbb{E}_{\mathbf{X},\theta_{1}}\left\{\hat{f}_{\operatorname{G}}(\mathbf{X};\mathcal{D}_{n},\theta_{1})^{2}\mathbb{1}\left\{J(\mathbf{x};\mathcal{D}_{n},\theta_{1})\cap T\neq\emptyset\right\}\right\}\\ &\leq M^{2}\mathbb{P}_{\mathbf{X},\theta_{1}}\left\{J(\mathbf{X};\mathcal{D}_{n},\theta_{1})\cap T\neq\emptyset\right\}.\end{split}$ (J.12)
@@
$\mathbb{E}_{\mathcal{D}_{n}}\left\{\mathbb{P}_{\mathbf{X},\theta_{1}}\left\{J(\mathbf{X};\mathcal{D}_{n},\theta_{1})\cap T\neq\emptyset\right\}\right\}\leq\delta,$ (J.13)
@@
$\mathbb{P}_{\mathbf{X},\theta_{1}}\left\{J(\mathbf{x};\mathcal{D}_{n},\theta_{1})\cap T\neq\emptyset\right\}\leq 2\delta.$ (J.14)
@@
$\mathbb{E}_{\Theta}\left\{R\left(\hat{f}_{\operatorname{G-E}}(-;\mathcal{D}_{n},\Theta)\right)\right\}\geq\frac{\operatorname{Var}\left\{r_{\textsf{MSP}}(\mathbf{X})\right\}}{4}.$ (J.15)
@@
Appendix K Comparisons with Neural Networks Trained by SGD
@@
[…]
for any $i\in\{1,2,\ldots,r\}$. As can easily be seen, MSP functions are equivalent to leap-$1$ functions, while non-MSP functions are categorized as leap-$2$, leap-33, and so on. A Boolean monomial of degree $k$ (i.e., a $k$-parity) is an example of a leap-$k$ function. With this definition, Abbe et al. (2023) (see Conjecture 1 therein) conjectured that the time and sample complexity required to learn a leap-$k$ function using online SGD on a two-layer NN is $\tilde{\Theta}\left(d^{\max\left\{k-1,1\right\}}\right)$.
If this conjecture is true, then combining it with the results in our paper yields rigorous theoretical performance gaps between two-layer NNs trained with online SGD (henceforth denoted as 2LNNOSGD) and greedy regression trees and ensembles: More specifically, for leap-$1$ functions, that is, those satisfying MSP, CART requires a sample complexity of just $O(2^{s}\log d)$ which is much smaller than the $\tilde{\Theta}(d)$ sample complexity required by 2LNNOSGD. On the other hand, for leap-$k$ functions for any $k>1$, CART requires a sample complexity of $\exp(\Omega(d))$, which is much larger than the $\tilde{\Theta}\left(d^{k-1}\right)$ sample complexity required by 2LNNOSGD. Here we note that our lower bound in Theorem 1.1 holds even with noiseless observations and also with constant probability (see Theorem J.1), so that both classes of learning algorithms are put on exactly the same theoretical footing.
@@
The lower bounds portion of Abbe et al. (2023)’s conjecture is suggested by existing Statistical Query (SQ) and Correlational Statistical Query (CSQ) lower bounds. SQ and CSQ are generic models of constrained computation (Kearns, 1998; Feldman et al., 2021). In the SQ model, the analyst has access to a sequence of queries, each of which returns an approximation of the expectation $\mathbb{E}\left\{\psi(\textbf{X},Y)\right\}$ for some real-valued function $\psi$ that may change from query to query. The CSQ model further constrains the queries to be of the form $\mathbb{E}\left\{\phi(\textbf{X})Y\right\}$ for some real-valued function $\phi$. Under the SQ model, it is known that $\Omega\left(d^{k}\right))$ queries are required to learn $k$-parities (Barak et al., 2022), while Abbe et al. (2023) proved a CSQ lower bound of the same order. Now, Abbe et al. (2021b) showed that SGD is no more powerful than the SQ model if the stochastic gradients have limited precision, but it is unclear to us whether this holds for the SGD regime in 2LNNOSGD. Assuming the reduction holds, then the SQ and CSQ lower bounds translate to a lower sample complexity bound of $\tilde{\Theta}\left(d^{k-1}\right)$ (Kou et al., 2024).
@@
Appendix L Additional empirical results
Figure 3: Expected MSE $R(\hat{f}_{\operatorname{CART}},f)$ for CART fitted to $Y=X_{1}X_{2}+\alpha X_{1}+\epsilon$, with $\alpha=0$ (left panel) and $\alpha=0.02$ (right panel) when $\epsilon\sim N(0,0.01)$. CART is fitted with a minimum impurity decrease stopping criterion.
Figure 4: Split coverage for features $X_{2}$ and $X_{3}$ for CART fitted to $Y=X_{1}X_{2}+\epsilon$ with $\epsilon\sim N(0,0.01)$, without any stopping criterion. Split coverage is defined in (57).
Figure 5: Split coverage for features $X_{2}$ and $X_{3}$ for CART fitted to $Y=X_{1}X_{2}+\alpha X_{1}$ with $\alpha=0.02$, without any stopping criterion. Split coverage is defined in (57).
Figure 6: Split coverage for features $X_{2}$ and $X_{3}$ for CART fitted to $Y=X_{1}X_{2}+\alpha X_{1}+\epsilon$ with $\alpha=0.02$ and $\epsilon\sim N(0,0.01)$, without any stopping criterion. Split coverage is defined in (57).